OpenStudy (lgbasallote):

a thermometer reading 70 F is taken outside. five minutes later the thermometer reads 40 F. thirty five minutes after being taken outside,the thermometer reading is within one half a degree of the outside temperature find the value of the outside temperature assuming it is constant

OpenStudy (lgbasallote):

given @t = 0 T = 70 @t = 5 T = 40 @t = 35 T = \(T_m) - 0.5?

OpenStudy (lgbasallote):

hmm maybe... \[\Large \ln(\frac{40 - T_m}{60 - T_m}) = 4k\]

OpenStudy (lgbasallote):

what does "one half a degree of the outside temperature" mean??

mathslover (mathslover):

may be that it means 1/2 of the atmospheric temp

OpenStudy (lgbasallote):

i remember something like Tm - 0.5 and Tm + 0.5 i dont remember exactly

mathslover (mathslover):

ok ! so u want the soln of this question right?

OpenStudy (lgbasallote):

of course :)

mathslover (mathslover):

So u r studying the Newtons' laws of cooling right ?

ganeshie8 (ganeshie8):